12 Question ( point) a See page 4 Nitrogen forms a number of compounds with oxygen, for example nitrous oxide, N2O, is 63.65% N by mass. 3rd attempt See Periodic Table See Hint A second compound, which contains two atoms of nitrogen per molecule, is 25.94% N by mass. How many oxygen atoms are in the second compound? Write your answer as a whole number.
Answer #1

Mass of two nitrogen atoms = 14×2 = 28 g

(Atomic mass of N is 14g)

Molar mass of compound = 28 + 16n

(n is number of oxygen atoms)

So, Percent of nitrogen = 28×100/(28+16n) = 25.94

(28+16n) = 2800/25.94

28+16n = 107.94

16n = 107.94-28

n = 79.94/16 = 5

So, 5 oxygen atoms are present in second compound.

Please give feedback!